# Which Demographics do LLMs Default to During Annotation?

### v1

@@ Abstract
Demographics and cultural background of annotators influence the labels they assign in text annotation – for instance, an elderly woman might find it offensive to read a message addressed to a “bro”, but a male teenager might find it appropriate. It is therefore important to acknowledge label variations to not underrepresent members of a society. Two research directions developed out of this observation in the context of using large language models (LLM) for data annotations, namely (1) studying biases and inherent knowledge of LLMs and (2) injecting diversity in the output by manipulating the prompt with demographic information. We combine these two strands of research and ask the question to which demographics an LLM resorts to when no demographics is given. To answer this question, we evaluate which attributes of human annotators LLMs inherently mimic. Furthermore, we compare non-demographic conditioned prompts and placebo-conditioned prompts (e.g., “you are an annotator who lives in house number 5”) to demographics-conditioned prompts (“You are a 45 year old man and an expert on politeness annotation. How do you rate {instance}”). We study these questions for politeness and offensiveness annotations on the Popquorn data set, a corpus created in a controlled manner to investigate human label variations based on demographics which has not been used for LLM-based analyses so far. We observe notable influences related to gender, race, and age in demographic prompting, which contrasts with previous studies that found no such effects.
[…]
booktabs
@@
Christopher Bagdon, Aidan Combs, Lynn Greschner, Roman Klinger, Jiahui Li, Sean Papay, Nadine Probol, Yarik Menchaca Resendiz, Johannes Schäfer, Aswathy Velutharambath, Sabine Weber,  and Amelie Wührlthanks: All authors contributed equally. Fundamentals of Natural Language Processing, University of Bamberg {firstname.lastname@uni-bamberg.de}
@@ 1 Introduction
With the advent of (instruction-tuned) large language models, automatic data annotation and zero or few shot predictions became more popular (Brown et al., 2020). However, language models do not provide the same diversity as human annotators do by simply prompting them in a zero-shot setup, which can lead to a lower performance of the labeling process (Bagdon et al., 2024). To mitigate this problem, it is important to understand whether large language models exhibit a default persona when acting as annotators and how this can be controlled to ensure greater diversity in annotations. One idea to address this limitation is socio-demographic prompting (Muscato et al., 2024), where we specify characteristics like age, gender, ethnicity or socio-economic status.
Previous research has explored the effects of using prompts informed by demographic or cultural contexts on model predictions, but did not find consistent patterns (Beck et al., 2024; Mukherjee et al., 2024). We build on top of this work and particularly contribute in two directions of socio-demographic prompting in a large demographic-annotated data set, namely Popquorn (Pei and Jurgens, 2023) that did not undergo an analysis from the LLM-perspective yet. (1) We study if large language models default to a particular demographic, i.e., we evaluate if the LLM-based predictions default to a particular demographic when not being conditioned. (2) Further, we use the idea of placebos (i.e., irrelevant information that should not change the output of the model) and compare demographic prompts to evaluate their robustness. More concretely, we answer the following questions.
@@
    How do task properties of offensiveness rating vs. politeness rating influence the role of demographic information or are observed patterns consistent across the two application tasks?
@@
    Are observed patterns consistent across different models?
@@ 2.1 Inherent Knowledge and Biases in LLMs
Large language models may be understood as databases that store information encoded in the training data (Petroni et al., 2019). These data is stored in a probabilistic manner, which allows these models to show sometimes unexpected generalization capabilities beyond the original instructions available in the training data, i.e., emergent abilities (Wei et al., 2022). It has further been shown that prompting may show properties of adapting parameters through backpropagation purely during inference (von Oswald et al., 2023). Therefore, it is reasonable to consider prompting even as a programming task (Beurer-Kellner et al., 2023).
@@
In cases, in which subjective properties are requested, the models therefore rely on “emergent” abilities. There is some research in understanding the “stances” or “worldviews” encoded in language models. For instance, Motoki et al. (2023); Feng et al. (2023) study political biases and how these lead to unfair models. Ceron et al. (2024) study the reliability and robustness of such stances. Wright et al. (2024) extend such analysis to more fine-grained opinions. The opinions expressed by models correspond more to particular subsets of a society than to others (Santurkar et al., 2023).
@@ 2.2 Perspectivism
Disagreement is related to confidence of annotators (and models), an aspect that received some attention. Baumler et al. (2023) made use of this property in the context of active learning, to understand which instances require annotations by multiple people, to understand the disagreements. Troiano et al. (2021); Baan et al. (2024) showed that annotator’s own confidence predicts inter-annotator agreement scores.
To understand label disagreement better, an increasing number of corpora is published with more detailed information on the annotators. Troiano et al. (2023) annotated event reports for emotions and appraisals and collected demographic information, personality, current emotional state of both the person who lived through the event and multiple annotators that read the event description. Plepi et al. (2022) studied the role of demographics, automatically extracted from the data, on the perception of social norms. Bizzoni et al. (2022) discusses the role of individual differences on the judgement of literary quality. Romberg (2022) integrates perspectivism in argument mining, by making explicit the subjective nature of argument interpretation. May et al. (2024) study the role of numbers in social judgements, under consideration of demographics. Frenda et al. (2023) study the perception of irony, depending on nationality, employment status, student status, ethnicity, age, and gender. Sachdeva et al. (2022) measures aspects of hate speech, including sentiment, disrespect, insult, attacking/defending, humiliation, inferior/superior status, dehumanization, violence, genocide, and a 3-valued hate speech benchmark label. They study these variables under the condition of identity target groups and annotator demographics. Xu et al. (2023) look into disagreement of legal case outcome differences. Next to subjective tasks, perspectivism has also been considered in seemingly objective tasks, for instance in named entity recognition (Peng et al., 2024) and natural language inference (Gruber et al., 2024).
The large Popquorn corpus (Pei and Jurgens, 2023) has been created to particularly study perspectivism of annotators and consider the tasks of offensiveness detection, question answering, text rewriting and style transfer, and politeness rating, depending on demographic variables. We use this corpus because it has been created for the study of perspectivism, but it did not undergo analyses from the perspective of large language models.
@@ 2.4 Socio-Demographic Prompting
The automatic annotation with language models does not correspond to human annotation – humans have previous world knowledge, experiences, and perspectives on a matter that differ individually. This comes with issues of automatic annotation. For instance, Bagdon et al. (2024) showed that the diversity of annotations that is beneficial in comparative annotations by multiple people is not straight-forwardly replaced by multiple runs of a language model. Lee et al. (2023) focused on this aspect and studied the alignment of the distribution of labels with human annotataed labels on natural language inference tasks.
Most relevant for our work are the following studies. Beck et al. (2024) study the impact of demographic information on subjective annotation tasks. They find that model variation is larger across other parameters than demographic information (e.g., age and gender) in the prompt. Mukherjee et al. (2024) study a similar setup, namely for cultural aspects of the annotators (e.g., food preferences). To understand if the variables influence the annotation in a systematic way, they compare the predictions to what they call “placebos” – information that should not influence the prediction but looks like relevant parameters (favorite planet or house number). Finally, there is a set of studies that investigate biases on large language models (Cheng et al., 2023; Santy et al., 2023, i.a.).
@@ 3.1 Data Sets
The Popquorn data set consists of 45,000 annotations from 1,484 annotators and includes information on the annotators on gender, race, age, occupation and education. It is partioned into multiple NLP tasks, from which we use the task on offensiveness ratings and politeness ratings. To equalize the number of annotations per instance, we sample three annotations for each instance. When sampling, we omit the annotations where the annotator preferred not to disclose demographic information. We exclude a data instance because it has fewer than three annotations remaining. The two subsets of the sample which we use in our experiments consist of 4,500 annotations of 1,500 instances for offensiveness ratings and 11,151 annotations of 3,717 instances for politeness ratings.
@@ 3.2 Prompt Setup
We use three different prompt setups, namely socio-demographically informed (SD) prompts, placebo-informed (P) prompts, and prompts without additional information (N prompts). We show the templates for these prompts in Table 1. The input to the LLM consist of a concatenation of the system message with the respective text to be classified. The values for demographic attributes are sampled from the demographics of the annotators. The sets of values for the placebo attributes are displayed in Table 2. During prompting, each placebo attribute value is randomly sampled from the respective set.
@@ 3.3 Model Choice and Access
We use two LLMs, namely GPT-4o [^1] https://openai.com/index/hello-gpt-4o/ and Claude [^2] https://www.anthropic.com/news/claude-3-5-sonnet, claude-3-5-sonnet-20240620. We access both models via their respective APIs. The total cost for using GPT-4o amounted to \$54, while the cost for using Claude was \$60.
@@ 3.4 LLM Output Parsing
There were few cases where the output could not be parsed. We encountered 22 such cases with GPT-4o and three with Claude. Of the failed cases with GPT-4o, 14 were due to the input text being in Polish instead of English. This resulted in the model predicting labels in Polish, despite the prompt specifying an English-only label set. The remaining eight cases involved very short text samples or questions, which the model misidentified and prompted it to request clarification instead of performing the classification task.
For Claude, the three instances of failure involved actor names in the text. This triggered the model’s copyright protection protocols, preventing it from reproducing or paraphrasing potentially copyrighted content. Given the total number of instances analyzed in our experiments, these numbers are relatively insignificant. Consequently, we disregarded the output in these cases and assigned the labels “not offensive” or “not polite”.
@@ 3.5 Evaluation Settings
We conduct various analyses of the labels generated by the LLMs, utilizing the different prompt setups described in Section 3.2. First, we evaluate the extent to which the LLMs replicate human annotations by comparing the predictions made with non-demographic (N) prompts to the gold standard annotations. Second, we examine the impact of socio-demographic prompting by comparing the models’ automatic annotations produced with SD prompts to those produced with N prompts. In this analysis, we assess the differences in the LLMs’ annotations when demographic data is included versus when it is omitted. Thirdly, we investigate the effect of placebo prompting in a similar manner by comparing the automatic annotations generated with P prompts to those created with N prompts.
@@ 4 Results
In this section, we present the results of our experiments based on the annotations generated by the LLMs for two labeling tasks. Each of our research questions is addressed with particular results derived from the various experiments conducted. Detailed results for individual prompts given in Section A.2 show that the two corresponding prompts for no-info promping behave similarly. We also observed this for the other prompt types (see Table 1). Thus, we here report all results as an average over the two respective prompts for the different prompt types.
@@
Table 3: Coefficients indicating the effect of particular human demographic categories on the distance between human and LLM annotations. Statistical significance is calculated based on standard error (see Section A.1) and is here marked by asterisks: * corresponds to P $\leq$ 0.05, ** to P $\leq$ 0.01, and *** to P $\leq$ 0.001.
@@
Which socio-demographic attributes of human annotators does an LLM inherently mimic in the absence of explicit information (N prompts)? Table 3 shows the results for this analysis which corresponds to the first setting described in Section 3.5. We combine socio-demographic attributes that are represented in only a few cases. The table specifies the reference categories (ref.) used for each of the categorical variables. We define such reference to report coefficients that describe the relation to other socio-demographic categories.
@@
Table 3 shows coefficients for the two models (GPT-4o and Claude) and two classification tasks. The coefficients can be interpreted as the effect of particular human demographic categories on the distance between human and LLM annotations. The dependent variables to compute the scores are the difference between the absolute values of the human annotation and the annotation of the respective LLM with non-demographic prompting. The independent variables are the human annotator demographic attribute values. Positive coefficients indicate that the LLM is further from the human annotators in that demographic category, i.e., less accurate. So, categories with positive, statistically significant coefficients are those categories which the respective model does not default to.
@@
This analysis (see Table 3) reveals significant biases in the LLMs’ predictions based on demographic factors, particularly concerning race and age. Specifically, the models demonstrate a tendency to align more closely with annotations provided by persons identying as White as opposed to those with the value Black or African American. The distance between their annotations measures .19 to .22 Likert scale points for offensiveness rating and .14 to .15 points for politeness rating. Additionally, the LLMs are progressively less accurate in reflecting the views of older individuals at offensiveness rating, with a .01 point increase in distance for each year of age. In terms of educational background, the models exhibit a greater discrepancy from those with less than a high school education compared to those with higher educational levels, ranging from .34 to .57 points for politeness rating. These patterns suggest that the models’ training data may lack representation for these demographic groups, affecting its annotation accuracy.
@@
Notably, the analysis found no significant effects related to gender or occupational status. Overall, while the statistically significant differences in annotation distances suggest some demographic biases, these differences are modest compared to the full Likert scale range. This complexity makes it challenging to establish a clear default persona for the LLMs. Therefore, we perform a second analysis in the following.
@@
Table 4: Sample sizes and mean distance scores of demographic-prompting predictions in comparison to no-info-prompting predictions for models GPT-4o and Claude at two rating tasks. Variables with very few cases are shown in gray to indicate a lack or reliability of these numbers.
The Effect of Demographic Prompting. In a second approach, we examine how the inclusion of demographic information in the prompt (cf. second setting in Section 3.5) influences the automatic annotation outputs of the models. We compare SD prompts with N prompts. This allows us to infer the demographics that the models defaults to.
@@
Claude shows a notable pattern for the politeness task concerning the age socio-demographic attribute. With an increasing age, the prediction difference increases (from 16 to .26). Similarly, the occupation attribute reflects the highest difference for the value Retired. This indicates that when Claude is prompted to act as an older individual, it exhibits an increased sensitivity to politeness.
@@
We evaluate if these results are as pronounced as the observations based on Table 4 as described above. Overall, the results of the placebo prompting (P prompts) in comparison to (N prompts) do not reveal any notable differences for specific attribute values. The scores remain consistently stable across these comparisons. Therefore, we show the results only in the appendix, in Section A.3. Consequently, we conclude that the changes in model behavior are indeed more pronounced for socio-demographic prompting compared to the inclusion of non-relevant additional information.
4.3 RQ3: How do task properties of offensiveness rating vs. politeness rating influence the role of demographic information or are observed patterns consistent across the two application tasks?
We investigate how task properties influence the role of demographic information in prompting and whether these patterns remain consistent across different application tasks. Our general conclusion indicates that there is no clearly distinguished default persona of the models for both tasks. Table 3 suggest that the results are in general consistent for the two tasks. However, we did also highlight some task-specific tendencies above at the analysis regarding RQ1. Notably, there are also differences in the absolute values of the prediction distances when comparing demographic prompting to no-information prompting across the two tasks. The average prediction difference associated with socio-demographic prompting for offensiveness rating is calculated to be 0.19. In contrast, the average for politeness rating is substantially higher at .25. Similar results are evident for the model Claude (0.16 for offensiveness rating and 0.21 for politeness rating). This suggests that, in general, the LLMs are more influenced by demographic information at rating politeness than rating offensiveness.
4.4 RQ4: Are observed patterns consistent across different models?

### v2

@@ Abstract
Demographics and cultural background of annotators influence the labels they assign in text annotation – for instance, an elderly woman might find it offensive to read a message addressed to a “bro”, but a male teenager might find it appropriate. It is therefore important to acknowledge label variations to not underrepresent members of a society. Two research directions developed out of this observation in the context of using large language models (LLM) for data annotations, namely (1) studying biases and inherent knowledge of LLMs and (2) injecting diversity in the output by manipulating the prompt with demographic information. We combine these two strands of research and ask the question to which demographics an LLM resorts when no demographics is given. To answer this question, we evaluate which attributes of human annotators LLMs inherently mimic. Furthermore, we compare non-demographic conditioned prompts and placebo-conditioned prompts (e.g., “you are an annotator who lives in house number 5”) to demographics-conditioned prompts (“You are a 45 year old man and an expert on politeness annotation. How do you rate {instance}”). We study these questions for politeness and offensiveness annotations on the Popquorn data set, a corpus created in a controlled manner to investigate human label variations based on demographics which has not been used for LLM-based analyses so far. We observe notable influences related to gender, race, and age in demographic prompting, which contrasts with previous studies that found no such effects.
@@
Johannes Schäfer, Aidan Combs, Christopher Bagdon, Jiahui Li, Nadine Probol, Lynn Greschner, Sean Papay, Yarik Menchaca Resendiz, Aswathy Velutharambath, Amelie Wührlthanks: These authors contributed equally., Sabine Weber,  and Roman Klinger Fundamentals of Natural Language Processing, University of Bamberg roman.klinger@uni-bamberg.de
@@ 1 Introduction
With the advent of (instruction-tuned) large language models, automatic data annotation and zero or few shot predictions became more popular (Brown et al., 2020). However, language models do not provide the same diversity as human annotators do by simply prompting them in a zero-shot setup, which can lead to a lower performance of the labeling process (Bagdon et al., 2024). To mitigate this problem, it is important to understand whether large language models exhibit a default persona when acting as annotators and how this can be controlled to ensure greater diversity in annotations. One concern is that large language models may reflect biases present in their training data, which can disproportionately emphasize certain viewpoints. This raises the issue of potential marginalization of minority perspectives in the outputs generated by these models, making it crucial to address this bias for equitable representation in annotation tasks. One idea to address this limitation is socio-demographic prompting (Muscato et al., 2024), where we specify characteristics like age, gender, ethnicity or socio-economic status.
Previous research has explored the effects of using prompts informed by demographic or cultural contexts on model predictions, but did not find consistent patterns (Beck et al., 2024; Mukherjee et al., 2024). We build on top of this work and particularly contribute in two directions of socio-demographic prompting in a large demographic-annotated data set, namely Popquorn (Pei and Jurgens, 2023) that has not undergone an analysis from the LLM-perspective yet. (1) We study if large language models default to a particular demographic, i.e., we evaluate if the LLM-based predictions more closely align with those of people of a particular demographic when not being conditioned. (2) Additionally, we utilize placebos as suggested by Mukherjee et al. (2024) – irrelevant information that should not affect the model’s output – to compare the impact of demographic prompts and evaluate the model’s prediction stability. More concretely, we answer the following questions.
@@
    How do task properties of offensiveness rating vs. politeness rating influence the role of demographic information in prompts?
@@
    Are observed patterns consistent across different large language models?
@@ 2.1 Inherent Knowledge and Biases in LLMs
Large language models may be understood as databases that store information encoded in the training data (Petroni et al., 2019). These data is stored in a probabilistic manner, which allows these models to show sometimes unexpected generalization capabilities beyond the original instructions available in the training data, i.e., emergent abilities (Wei et al., 2022). It has also been shown that prompting can adapt parameters during inference, similar to how backpropagation works (von Oswald et al., 2023). Therefore, it is reasonable to view prompting as a form of programming, as it involves manipulating model behavior by crafting specific inputs to achieve desired outputs (Beurer-Kellner et al., 2023).
@@
In cases in which subjective properties are requested, the models therefore rely on “emergent” abilities. There is some research in understanding the “stances” or “worldviews” encoded in language models. For instance, Motoki et al. (2023); Feng et al. (2023) study political biases and how these lead to unfair models. Ceron et al. (2024) study the reliability and robustness of such stances. Wright et al. (2024) extend such analysis to more fine-grained opinions. The opinions expressed by models correspond more to particular subsets of a society than to others (Santurkar et al., 2023).
@@ 2.2 Perspectivism
Disagreement is related to confidence of annotators (and models), an aspect that has received some attention. Baumler et al. (2023) made use of this property in the context of active learning, to understand which instances require annotations by multiple people in order to to understand the disagreements. Troiano et al. (2021); Baan et al. (2024) showed that annotator’s own confidence predicts inter-annotator agreement scores.
To understand disagreement in labeling better, more and more corpora are being published with more detailed information about the annotators. Troiano et al. (2023) annotated event reports for emotions and appraisals and collected demographic information, personality, and current emotional state of both the person who lived through the event and multiple annotators that read the event description. Plepi et al. (2022) studied the role of demographics, automatically extracted from the data, on the perception of social norms. Bizzoni et al. (2022) discuss the role of individual differences on the judgement of literary quality. Romberg (2022) integrates perspectivism in argument mining, by making explicit the subjective nature of argument interpretation. May et al. (2024) study the effect of demographics on the role of numbers in social judgements. Frenda et al. (2023) study how the perception of irony varies by nationality, employment status, student status, ethnicity, age, and gender. Sachdeva et al. (2022) measures aspects of hate speech, including sentiment, disrespect, insult, attacking/defending, humiliation, inferior/superior status, dehumanization, violence, genocide, and a 3-valued hate speech benchmark label. They study these variables under the condition of identity target groups and annotator demographics. Xu et al. (2023) look into disagreement of legal case outcome differences. Next to subjective tasks, perspectivism has also been considered in seemingly objective tasks, for instance in named entity recognition (Peng et al., 2024) and natural language inference (Gruber et al., 2024).
The large Popquorn corpus (Pei and Jurgens, 2023) has been created specifically to study perspectivism of annotators and the relationship between demographics and annotations in the tasks of offensiveness detection, question answering, text rewriting and style transfer, and politeness rating. We use this corpus because it has been created for the study of perspectivism, but it has not yet been used to analyze large language models.
@@ 2.4 Socio-Demographic Prompting
Automatic annotation with language models is not a replacement for human annotation. Humans have previous world knowledge, experiences, and perspectives on a matter that differ individually. LLMs have difficulties replicating these differences, causing issues in annotation. For instance, Bagdon et al. (2024) showed that the diversity of annotations that is beneficial in comparative annotations by multiple people is not straight-forwardly replaced by multiple runs of a language model. Lee et al. (2023) focused on this aspect and found poor alignment of the distribution of labels predicted by LLMs with human annotations on natural language inference tasks.
Most relevant for our work are the following studies. Beck et al. (2024) study the impact of demographic information on subjective annotation tasks. They find that model variation is larger across other parameters such as prompt formulation techniques than demographic information (e.g., age and gender) in the prompt. Mukherjee et al. (2024) examine cultural aspects like food preferences and find that most language models exhibit significant response variability, casting doubt on the reliability of socio-demographic prompting. To understand if the variables influence the annotation in a systematic way, they compare the predictions to what they call “placebos” – information that should not influence the prediction but looks like relevant parameters (favorite planet or house number).
Finally, there is a set of studies that investigate biases on large language models (Cheng et al., 2023; Santy et al., 2023, i.a.). Their findings suggest that these models may generate outputs reflecting racial stereotypes and exhibit some performance disparities across different demographic groups, indicating potential biases in their design and outputs.
@@ 3.1 Data Sets
The Popquorn data set consists of 45,000 annotations from 1,484 annotators and includes information on annotators’ gender, race, age, occupation and education. It is partioned into multiple NLP tasks. We use data from the offensiveness and politeness rating tasks. To equalize the number of annotations per instance, we sample three annotations for each instance. When sampling, we omit the annotations where the annotator preferred not to disclose demographic information. We exclude data instances with fewer than three annotations remaining. The two subsets of the sample which we use in our experiments consist of 4,500 annotations of 1,500 instances for offensiveness ratings and 11,151 annotations of 3,717 instances for politeness ratings.
@@ 3.2 Prompt Setup
We use three different prompt setups, namely socio-demographically informed (SD) prompts, placebo-informed (P) prompts, and prompts without additional information (N prompts). We show the templates for these prompts in Table 1. The input to the LLM consist of a concatenation of the system message with the respective text to be classified. The values for the demographic attributes are taken from the demographic data of the annotators included in the sample. The sets of values for the placebo attributes are displayed in Table 2. During prompting, each placebo attribute value is randomly sampled from the respective set.
@@ 3.3 Model Choice and Access
We use two LLMs, namely GPT-4o [^1] https://openai.com/index/hello-gpt-4o/ and Claude [^2] https://www.anthropic.com/news/claude-3-5-sonnet, claude-3-5-sonnet-20240620. We access both models via their respective APIs. The cost for using GPT-4o was \$54 and the cost for using Claude was \$60.
@@ 3.4 LLM Output Parsing
There were few cases where the output could not be parsed. We encountered 22 such cases with GPT-4o and three with Claude. Of the failed cases with GPT-4o, 14 were due to the input text from the original data set being in Polish instead of English. This resulted in the model predicting labels in Polish, despite the prompt specifying an English-only label set. The remaining eight cases involved very short text instances or questions, which the model misinterpreted. In these cases, GPT-4o asked for further input instead of performing the classification task.
For Claude, the three instances of failure involved actor names in the text. This triggered the model’s copyright protection protocols, preventing it from reproducing or paraphrasing potentially copyrighted content. Given the total number of instances analyzed in our experiments, this failure rate is insubstantial. Consequently, we disregarded the output in these cases and assigned the labels “not offensive” or “not polite”.
@@ 3.5 Evaluation Settings
We conduct various analyses of the labels generated by the LLMs, utilizing the different prompt setups described in Section 3.2. First, we examine the extent to which LLM predictions align with the judgments of different human annotators by comparing the outputs generated from non-demographic (N) prompts to the multiple annotations available for each instance. Second, we examine the impact of socio-demographic prompting by comparing the models’ automatic annotations produced with SD prompts to those produced with N prompts. In this analysis, we assess the differences in the LLMs’ annotations when demographic data is included versus when it is omitted. Thirdly, we investigate the effect of placebo prompting in a similar manner by comparing the automatic annotations generated with P prompts to those created with N prompts.
@@ 4 Results
In this section, we present the results of our experiments based on the annotations generated by the LLMs for two labeling tasks. Each of our research questions is addressed with particular results derived from the various experiments conducted. Detailed results for individual prompts given in Section A.2 show that the two N prompts behave similarly. We also observed this for the other prompt types (see Table 1). Thus, we here report all results as an average over the two respective prompts for the different prompt types.
@@
Table 3: Coefficients indicating the effect of particular human demographic categories on the distance between human and LLM annotations, calculated using mixed-effects regression models with random intercepts for annotators and instances. Statistical significance is calculated using standard error (see Section A.1 for these values) and is here marked by asterisks: * corresponds to P $\leq$ 0.05, ** to P $\leq$ 0.01, and *** to P $\leq$ 0.001.
@@
Which socio-demographic attributes of human annotators does an LLM inherently mimic in the absence of explicit information (N prompts)? Table 3 shows the results for this analysis which corresponds to the first setting described in Section 3.5. We combine socio-demographic attributes that are represented in only a few cases. The table specifies the reference categories (ref.) used for each of the categorical variables.
@@
Table 3 shows the coefficients of regression models predicting absolute values of the distance between NLLM annotations and annotations provided by human annotators for the same instance. Independent variables are the sociodemographic characteristics of the human annotators. We report results separately for the two models (GPT-4o and Claude) and two classification tasks. The reported coefficients can be interpreted as the effect of particular human demographic categories on the distance between human and LLM annotations (change in distance between the category in question and the reference category). Positive coefficients indicate that the respective LLM is further from the human annotators in that demographic category than human annotators in the reference category, i.e., less accurate. So, categories with positive, statistically significant coefficients are those categories which the respective model does not default to.
@@
This analysis (see Table 3) reveals significant biases in the LLMs’ predictions based on demographic factors, particularly concerning race and age. Specifically, the models demonstrate a tendency to align more closely with annotations provided by persons identifying as White as opposed to those identifying as Black or African American. The distance between their annotations measures .19 to .22 Likert scale points for offensiveness rating and .14 to .15 points for politeness rating. Additionally, the LLMs are progressively less accurate in reflecting the views of older individuals at offensiveness rating, with a .01 point increase in distance for each year of age. In terms of educational background, the models exhibit a greater discrepancy from those with less than a high school education compared to those with higher educational levels, ranging from .34 to .57 points for politeness rating. One possible explanation for these discrepancies by annotator sociodemographics is that the models’ training data may lack representation for these demographic groups.
@@
Notably, the analysis found no significant effects related to gender or occupational status. This could be because the LLMs do not consistently favor any particular group of people on these dimensions, or it could be because there are in fact no systematic differences in human annotation along these dimensions. Overall, while the statistically significant differences in annotation distances suggest some demographic biases, we emphasize that these effects are small compared to the full Likert scale range. In addition, the R2 Marginal values for these models (reported in Appendix A.1) are quite low, indicating that the sociodemographic categories explain little of the variability in human–LLM annotation differences. This complexity makes it challenging to establish a clear default persona for the LLMs. Therefore, we perform a second analysis in the following.
@@
Table 4: Sample sizes and mean distance scores of demographic-prompting predictions in comparison to no-info-prompting predictions for models GPT-4o and Claude at two rating tasks. Variables with very few cases (count $\leq$ 100) are shown in gray to indicate a lack or reliability of these numbers.
The Effect of Demographic Prompting. In a second approach, we examine how the inclusion of demographic information in the prompt (cf. second setting described in Section 3.5) influences the automatic annotation outputs of the models. We compare SD prompts with N prompts. This allows us to infer the demographics that the models defaults to.
@@
Claude shows a notable pattern for the politeness task concerning the age socio-demographic attribute. With an increasing age, the prediction difference increases (from .16 to .26). Similarly, the occupation attribute reflects the highest difference for the value Retired. This indicates that when Claude is prompted to act as an older individual, it exhibits an increased sensitivity to politeness.
@@
We evaluate if these results are as pronounced as the observations based on Table 4 as described above. Overall, the results of the placebo prompting (P prompts) in comparison to N prompts do not reveal any notable differences for specific attribute values. The scores remain consistently stable across these comparisons. Therefore, we show the results only in the appendix, in Section A.3. Consequently, we conclude that the changes in model behavior are indeed more pronounced for socio-demographic prompting compared to the inclusion of non-relevant additional information.
4.3 RQ3: How do task properties of offensiveness rating vs. politeness rating influence the role of demographic information in prompts?
We investigate how task properties influence the role of demographic information in prompting and whether these patterns remain consistent across different application tasks. Our general conclusion indicates that there is no clearly distinguished default persona of the models for both tasks. Table 3 suggest that the results are in general consistent for the two tasks. However, we did also highlight some task-specific tendencies above at the analysis regarding RQ1. Notably, there are also differences in the absolute values of the prediction distances when comparing demographic prompting to no-information prompting across the two tasks. The average prediction difference associated with socio-demographic prompting of GPT-4o for offensiveness rating is calculated to be 0.19. In contrast, the average for politeness rating is substantially higher at 0.25. Similar results are evident for the model Claude (0.16 for offensiveness rating and 0.21 for politeness rating). This suggests that, in general, the LLMs are more influenced by demographic information at rating politeness than rating offensiveness.
4.4 RQ4: Are observed patterns consistent across different large language models?